OpenStudy (anonymous):

in word form. simply how does one multiply two powers of ten, and how do you divide two powers of ten?

OpenStudy (anonymous):

Can you explain that a bit more? I'm not sure if I understand the question - do you mean how does 2 to the power of 10 look like? Then again, this question can also just be outta my league O_O

OpenStudy (anonymous):

for multiplication you just add the powers e.g. 10^7 * 10^4 = 10^ 11 for dividing you subtract the power of the divisor from the power of dividend e.g. 10^7 ÷ 10^4 = 10^3